A nitrox ll gas mixture for scuba diving contains oxygen gas ar 53 atm and nitrogen gas at 94 atm. What is the total pressure, in torr?
Answer #1

According to Dalton's law of partial pressure , the total pressure exerted by a mixture of gases is equal to the sum of individual pressures of each gas.

1 atm = 760 torr.

Oxygen gas,

Pressure = 760 × 53

= 40280 torr

Nitrogen gas,

Pressure = 94 × 760

= 71440 torr

Total pressure = 71440 + 40280

= 111720 torr